A boat goes 200 miles downstream in the same time it can go 150 miles up stream. The speed of the current is 7 miles per hour. Find the speed of the boat in still water.
Answer by Paul(988)

Letting x be the speed of the boat
EQUATION:
200(x-7)=150(x+7)
200x-1400=150x+1050
50x=2450
x=49
Hence, the speed of the boat is 49mph.
Paul.